Fermi National Laboratory
 

Operation Summary for September 28, 2009 - September 30, 2009

 

Monday Sept. 28

The day shift began with Operations monitoring the Tevatron (TeV) store 7220 (85.08E30) and the Recycler stash (313.0E10), with the Antiproton Source stacking (19.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

Operations terminated store 7220 at 10 AM. TeV orbit studies followed until 2:30 PM.

Controls personnel replaced the crate controller for the TeV sector C3 crate.

Kicker personnel replaced the high voltage power supply for a MI kicker (KPS5S).

MI personnel conducted a study.

At 1:13 PM, a TeV RF station (TRF2) power amplifier (PA) developed a LCW leak. RF personnel had the PA repaired by 4 PM.

TeV personnel began a study.

At 5:12 PM, a TeV quadrupole power supply (C:BQ7) tripped off while changing polarities. EE Support technicians had it repaired early on the evening shift.

Operations put the TeV into shot setup at 4:47 PM.

Operations established store 7224 at 10:27 PM with an average initial luminosity (AIL) of 228.18E30.

Tuesday Sept. 29

The midnight shift began with Operations monitoring store 7224 (128.69E30) and the stash (256.6E10), with Pbar stacking (26.8E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 7224 at 4:12 AM.

Operations put the TeV into shot setup and established store 7225 at 5:50 AM with an AIL of 232.16E30.

TRF2 tripped off at 6:30 AM. It was left off.

Operations turned off the NuMI beam at 7:15 AM so NuMI could switch to anti-neutrino mode; it took most of the day shift to make the switch. A stacking-only timeline was loaded into the Time Line Generator. They resumed taking beam just before the end of the day shift.

MI personnel conducted a coalesced beam study.

Booster RF work caused some downtime during the day shift, as did the MI low-level RF.

The NuMI horn tripped off and caused some NuMI downtime.

The TeV A49 separator tripped off due to a load resistor over temperature fault. TeV personnel arrived, purged the system, and got the separator running again.

Operations terminated store 7225 at the end of the evening shift and put the TeV into shot setup.

Wednesday Sept. 30

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (355.4E10), with Pbar stacking (32.7E10), and with NuMI and MiniBooNE taking beam.

Operations established store 7226 at 2:09 AM with an AIL of 242.63E30.

At 7:04 AM, store 7226 aborted when a TeV quadrupole power supply (C:CQ7) tripped off. There was no quench. Investigation and repairs followed.

There was a short NuMI downtime when the horn supply tripped off.

The Plans for Wednesday

The plans for today are to stack, stash, and store. TeV personnel will repair C:CQ7. The Run Coordinator will allow mechanical technicians to search TeV sector E3 for an air leak. He will also allow P1 transfer line beam studies, Pbar water cage work, and, if possible, MI/Switchyard septa beam tuning.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/01/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory